# Characterizing the Hyperuniformity of Disordered Network Metamaterials

### v1

@@ Abstract
Advancements in materials design and manufacturing have allowed for the production of ordered and disordered metamaterials with diverse and novel properties. Hyperuniform two-phase heterogeneous materials, which anomalously suppress density fluctuations on large length scales compared to typical disordered systems, and network materials are two classes of metamaterial that have desirable physical properties. Recent focus has been placed on the design of disordered hyperuniform network metamaterials that inherit the desirable properties of both of these metamaterial classes. In this work, we focus on determining the extent to which network structures derived from the spatial tessellations of hyperuniform point patterns inherit the hyperuniformity of the progenitor point patterns. In particular, we examine the Delaunay, Voronoi, Delaunay-Centroidal, and Gabriel tessellations of nonhyperuniform and hyperuniform point patterns in two- and three-dimensional Euclidean space. We use the spectral density to characterize the density fluctuations of two-phase media created by thickening the edges of these tessellations in two dimensions and introduce a novel variance-based metric to characterize the network structures directly in two and three dimensions. We find that, while none of the tessellations completely inherit the hyperuniformity of the progenitor point pattern, the degree to which the hyperuniformity is inherited is sensitive to the tessellation scheme and the short- and long-range translational disorder in the point pattern, but not to the choice of beam shape when mapping the networks into two-phase media.
@@ I Introduction
Advancements in materials design and manufacturing have allowed for the production of ordered and disordered metamaterials with diverse and novel properties including, e.g., negative refractive indices [1], negative thermal expansion [2], or the ability to mimic the mechanical properties of biological tissues [3, 4, 5, 6]. Disordered hyperuniform materials are a particularly interesting class of metamaterial, in which infinite-scale density fluctuations are anomalously suppressed compared to those in ordinary disordered systems [7, 8]. Such materials have been demonstrated to have many desirable physical properties, such as complete isotropic photonic band gaps [9, 10, 11, 12] as well as optimal transport [13, 14] and mechanical properties [14, 15]. A second class of exciting metamaterials is network materials, whose low density and better strength-to-weight and stiffness-to-weight ratios [16] than bulk solids make them ideal for use in, e.g., medical and aerospace applications. Due to the advantages of these material classes, there have been recent efforts to design and fabricate disordered hyperuniform network metamaterials to capture the beneficial properties from both of these material classes [3, 6, 17, 12].
@@
Due to the great multitude of possible hyperuniform two-phase structures, there is interest in devising fast and simple methods for generating readily manufacturable hyperuniform structures. One possible avenue for doing so is the direct mapping of a hyperuniform point pattern to a two-phase material. While it is known that, e.g., the mapping of a hyperuniform point pattern to a packing of identical nonoverlapping spheres preserves the hyperuniformity of the progenitor point pattern [22], it is not generally known what mappings do or do not preserve hyperuniformity. Due to the desirable photonoic [23, 12], elastic, and conductive [14] properties of network structures derived from hyperuniform point patterns, there is a reasonable expectation that network structures inherit the hyperuniformity of their progenitor point patterns. Indeed, in Ref. 14, Torquato and Chen conjecture that the network structures derived from stealthy hyperuniform point patterns are also stealthy and hyperuniform.
@@
There has been a recent focus on characterizing the structures of networks generated via modifying existing point patterns in various ways. In Ref. [14], Torquato and Chen showed network materials derived from tessellations of disordered stealthy hyperuniform point patterns have near-optimal elastic and conductive physical properties. In addition, they demonstrated that different tessellation schemes are better for certain physical properties than others, however, the large-scale structural properties of these network materials were not directly assessed. Salvalaglio et al. [24] demonstrated that, by analyzing the topology of networks generated via persistent homology, one can diagnose the hyperuniformity of finite point patterns with a particular class of hyperuniform and nearly hyperuniform structure factors. Their method, however, does not examine the hyperuniformity of the networks generated with these methods. Moreover, Newby et al. [25] showed that one can use the Voronoi-area correlation function to determine if a Voronoi tessellation was generated from a hyperuniform, nonhyperuniform, or antihyperuniform point pattern. We also note here that Kim and Torquato [15] have devised a method to produce large disordered hyperuniform particle dispersions based on spatial tessellations of point patterns, but their procedure is distinct from what we consider herein.
@@
These works lack a direct characterization of the large-scale density fluctuations of the structures themselves derived from hyperuniform point patterns. In this work, for the first time, we characterize the degree to which the hyperuniformity of the point pattern is inherited by network structures composed of the edges of tiles from spatial tessellations of 2D and 3D hyperuniform point patterns (see Sec. II for more details). Specifically, we build on the configuration-to-network-to-3D printed structure pipeline introduced in Obrero et al. [17] for designing and manufacturing disordered network metamaterials. First, we generate uniformly randomized lattice (URL) [26] and disordered stealthy hyperuniform point patterns with varying degrees of short-scale translational disorder, as well as totally uncorrelated point patterns in $d=2,3$ under periodic boundary conditions. Then, we generate the Voronoi (V), Delaunay (D), Delaunay-Centroidal (C), and Gabriel (G) tessellations of these point patterns and use the edges of the tiles in these tessellations to form spatially embedded network structures.
@@
Subsequently, we characterize these structures in two ways. First, to characterize the structures of two-phase media derived from 2D networks across length scales, we apply a “thickness” to the network edges such that the network fills a non-zero fraction of the space, then compute the spectral density $\tilde{\chi}_{{}_{V}}(\mathbf{k})$. Second, inspired by traditional variance-based order metrics [27, 28, 29], we propose the use of a novel variance measure, the edge-length variance $\sigma^{2}_{\ell}(R)$, associated with the total edge length of a spatially embedded network within a spherical observation window of radius $R$, which we apply to our 2D and 3D networks. Such a variance measure allows us to analyze the structure of the networks themselves as opposed to two-phase structures derived from them, avoiding the need to make a particular choice for the edge thickness or shape when analyzing the structures.
@@
We find that the hyperuniformity of point patterns is not completely inherited by network structures derived from their spatial tessellations, and the degree to which hyperuniformity is inherited varies across different point patterns, tessellation types, and space dimensions. Such subtleties in the inheritance of hyperuniformity by the network structures are consistent with the varying rank-orderings of physical properties for the 2D networks derived from different tessellations of hyperuniform and nonhyperuniform point patterns in Ref. 14. More specifically, for 2D networks derived from disordered hyperuniform point patterns with small degrees of local translational order, we find that G networks do not inherit the hyperuniformity of the progenitor point patterns, while C networks tend to inherit the hyperuniformity most effectively. However, we find that the D, V, and C networks derived from these hyperuniform point patterns only exhibit hyperuniform density fluctuations up to a finite length scale, past which they exhibit nonhyperuniform density fluctuations. Despite not completely inheriting the hyperuniformity of the progenitor point patterns, D, C, and V networks with sufficiently small degrees of local translational disorder are effectively hyperuniform, and have values of the hyperuniformity index $H$ that are correlated to the degree of local translational disorder.
@@
We also demonstrate that our new proposed measure $\sigma^{2}_{\ell}(R)$ has analogous scaling behaviors to the number variance $\sigma_{N}^{2}(R)$, and picks up the same structural characteristics at intermediate- and large-length scales in our network structures as the well-established $\tilde{\chi}_{{}_{V}}(k)$ calculation. This indicates that the incomplete inheritance of hyperuniformity by the two-phase network media is intrinsic to the tessellation-based network structures and not a result of a particular choice of edge thickness (or shape, for $d=3$) when mapping the network structures to two-phase media. Having established the correspondence between traditional variance measures and $\sigma_{\ell}^{2}(R)$ for the 2D networks, we apply the same method to 3D networks and show that V networks derived from disordered hyperuniform point patterns tend to have the greatest suppression of large-scale density fluctuations, followed by C, D, and G networks. This change in optimal network type from C networks for $d=2$ to V networks for $d=3$ is not unexpected given that the currently known structures that minimize large-scale density fluctuations also change across space dimensions [18].
@@
The rest of the paper is organized as follows. Section II contains pertinent mathematical background for the methods used to generate and characterize the point patterns and networks used in this work. In Sec. III, we characterize the structures of these networks and propose a new variance-based network characterization method. In Sec. IV we offer concluding remarks and outlook for future studies.
@@ II Mathematical Background and Methods
In this section, we review the quantification of the pair statistics of point patterns and two-phase heterogeneous media and how they can be used to determine and characterize the degree of hyperuniformity in a given system. Then, we introduce the one nonhyperuniform and two hyperuniform point pattern models considered herein. Finally, we describe the four tessellations schemes we use to convert these point patterns into networks and the method used to convert these networks into two-phase media.
@@ II.1 Pair Statistics
$S(\mathbf{k})=1+\rho\tilde{h}(\mathbf{k})$ (1)
@@
$\tilde{\chi}_{{}_{V}}(\mathbf{k})=\frac{|\sum_{j=1}^{N_{p}}e^{-i\mathbf{k}\cdot{\mathbf{r}_{j}}}\tilde{m}(\mathbf{k};A)|^{2}}{V},\;(\mathbf{k}\neq 0)$ (6)
@@ II.3 Hyperuniform and Nonhyperuniform Model Point Patterns
In this work we consider one nonhyperuniform model and two hyperuniform model point patterns in $d=2,3$. The nonhyperuniform model we consider is a totally uncorrelated point process in $\mathbb{R}^{d}$. We generate these point patterns at a number density $\rho=1$ in a fixed $d$-dimensional box by uniformly randomly placing $N$ points independent of each other. Such a system with no spatial correlations has $g_{2}(r)=S(k)=1$ and thus is nonhyperuniform (i.e., $\alpha=0$) with $\sigma_{N}^{2}\sim R^{d}$ at large $R$.
One hyperuniform model we consider is the uniformly randomized lattice (URL) [26]. Here, we make the particular choice of displacing the points of the $\mathbb{Z}^{d}$ lattice by a random $d$-dimensional vector uniformly distributed on a scaled version of the unit cell of $\mathbb{Z}^{d}$, i.e., $a\mathcal{F}\equiv[-a/2,a/2)^{d}$, where $a$ controls the perturbation strength. Regardless of the exact value of $a>0$, the URL is a class I hyperuniform system with $\alpha=2$. One can tune the degree of local translational disorder in the point pattern with the specific choice of $a$, in particular, larger values of $a$ result in a greater degree of local translational disorder. Here, we consider values of $a$ between 0.1 and 1.75.
Additionally, we consider disordered stealthy hyperuniform point patterns, which have $S(k)=0$ for $0<k\leq K$ and thus are class I hyperuniform. We generate these point patterns using the collective coordinate optimization scheme starting from totally uncorrelated initial conditions [33]. The parameter $\chi$ is a dimensionless measure of the ratio of constrained degrees of freedom in the stealthy system (i.e., the number of wavenumbers less than $K$ constrained to be 0) to the total number of degrees of freedom (roughly $dN$, where $N$ is the number of particles in the system). Relatively unconstrained (low-$\chi$) disordered stealthy systems will have greater translational disorder on short length scales and as $\chi$ increases within the so-called “disordered regime” ($20<\chi<1/$ 2 for $d=2,3$ [19]) the degree of short length scale translational disorder decreases. In this work we consider $\chi=0.2,0.4,$ and 0.48.
@@ II.4 Network Generation
To generate the networks examined herein, we first map point patterns under periodic boundary conditions in $\mathbb{R}^{2}$ ($\mathbb{R}^{3}$) to sets of polygons (polyhedra) that tile the space using the Voronoi (V), Delaunay (D), Delaunay-centroidal (C), and Gabriel (G) tessellations (described below). Then, we produce a network from the edges that make up the boundary of each cell in the tessellation. Each polygonal cell in the Voronoi tessellation is the region of space closer to one point than to any other point in the progenitor point pattern. The Delaunay tessellation is the dual network of the Voronoi tessellation. The Delaunay-centroidal tessellation [9] is generated by connecting the centroid of each simplex in the Delaunay tesselation to the centroid of each neighboring simplex, i.e., those that share an edge (or face in $\mathbb{R}^{3}$). Finally, the Gabriel tessellation [34] is derived from the Delaunay tesselation by removing any edge if a sphere generated using that edge as its diameter contains another vertex in the tesselation.
@@
In addition to considering these sets of edges, in $\mathbb{R}^{2}$ we consider the two-phase medium produced when a rectangle of dimensions $2\delta\times b+2\delta$ is superposed on each edge, where $b$ is the Euclidean length of the original edge and $\delta$ is some prescribed amount by which the edge is “thickened.” Here, we choose $\delta$ to be equal to one twentieth of $\rho^{-1/d}$. We note here that the particular choice of beam thickness only impacts the small-scale structural characteristics of the medium and the large-scale density fluctuations are not significantly impacted. In Fig. 1 we show an example of each of these types of networks. Because this mapping of point patterns to two-phase media is a perturbation of the underlying network structure, we additionally desire a method to directly examine the network structure. We introduce and discuss the merits of such a method in Sec. III.2
III Results
III.1 Two-Dimensional Networks

### v2

@@ Abstract
Advancements in materials design and manufacturing have allowed for the production of ordered and disordered metamaterials with diverse properties. Hyperuniform two-phase heterogeneous materials, which anomalously suppress density fluctuations on large length scales compared to typical disordered systems, and network materials are two classes of metamaterials that have desirable physical properties. Recent focus has been placed on the design of disordered hyperuniform network metamaterials that inherit the desirable properties of both of these metamaterial classes. In this work, we focus on determining the extent to which network structures derived from the spatial tessellations of hyperuniform point patterns inherit the hyperuniformity of the progenitor point patterns. In particular, we examine the Delaunay, Voronoi, Delaunay-centroidal, and Gabriel tessellations of nonhyperuniform and hyperuniform point patterns in two- and three-dimensional Euclidean space. We use the spectral density to characterize the density fluctuations of two-phase media created by thickening the edges of these tessellations in two dimensions and introduce a variance-based measurement to characterize the network structures directly in two and three dimensions. We find that, while none of the tessellations completely inherit the hyperuniformity of the progenitor point pattern, the degree to which the hyperuniformity is inherited is sensitive to the tessellation scheme and the short- and long-range translational disorder in the point pattern, but not to the choice of beam shape when mapping the networks into two-phase media.
@@ I Introduction
Advancements in materials design and manufacturing have allowed for the production of ordered and disordered metamaterials with diverse properties including, e.g., negative refractive indices [1], negative thermal expansion [2], or the ability to mimic the mechanical properties of biological tissues [3, 4, 5, 6]. Disordered hyperuniform materials are a particularly interesting class of metamaterial, in which infinite-length scale density fluctuations are anomalously suppressed compared to those in ordinary disordered systems [7, 8]. Such materials have been demonstrated to have many desirable physical properties, such as complete isotropic photonic band gaps [9, 10, 11, 12] as well as optimal transport [13, 14] and mechanical properties [14, 15]. A second class of exciting metamaterials is network materials, whose low density and better strength-to-weight and stiffness-to-weight ratios [16] than bulk solids make them ideal for use in, e.g., medical and aerospace applications. Due to the advantages of these material classes, there have been recent efforts to design and fabricate disordered hyperuniform network metamaterials to capture the beneficial properties from both of these material classes [3, 6, 17, 12].
@@
Due to the great multitude of possible hyperuniform two-phase structures, there is interest in devising fast and simple methods for generating readily manufacturable hyperuniform structures. One possible avenue for doing so is the direct mapping of a hyperuniform point pattern to a two-phase material. While it is known that, e.g., the mapping of a hyperuniform point pattern to a packing of identical nonoverlapping spheres preserves the hyperuniformity of the progenitor point pattern [22], it is not generally known what mappings do or do not preserve hyperuniformity. Due to the desirable photonic [23, 12], elastic, and conductive [14] properties of network structures derived from hyperuniform point patterns, there is a reasonable expectation that network structures inherit the hyperuniformity of their progenitor point patterns. Indeed, in Ref. 14, Torquato and Chen suggest that such structures may be hyperuniform due to the presence of these desirable physical properties, but note that the exact conditions required to generate truly hyperuniform networks from disordered hyperuniform point patterns are yet to be identified.
@@
There has been a recent focus on characterizing the structures of networks generated via modifying existing point patterns in various ways. In Ref. [14], Torquato and Chen showed network materials derived from tessellations of disordered stealthy hyperuniform point patterns have near-optimal elastic and conductive physical properties. In addition, they demonstrated that different tessellation schemes are better for certain physical properties than others, however, the large-scale structural properties of these network materials were not directly assessed. Similarly, Raj et al. [24] examine the transport properties of networks derived from the tessellations of hyperuniform and nonhyperuniform point patterns, as well as their local geometric properties. Moreover, in Refs. 25 and 26, the authors show that disordered network structures generated by adding topological defects to ordered two-dimensional networks have hyperuniform vertex locations and unique electronic transport properties. Salvalaglio et al. [27] demonstrated that, by analyzing the topology of networks generated via persistent homology, one can diagnose the hyperuniformity of finite point patterns with a particular class of hyperuniform and nearly hyperuniform structure factors. Their method, however, does not examine the hyperuniformity of the networks generated with these methods. Moreover, Newby et al. have characterized the local fluctuations in Voronoi cell volumes [28] and density fluctuations in the nodes of Delaunay networks [29] derived from hyperuniform and nonhyperuniform point patterns using several correlation functions and network characterization methods. We also note here that Kim and Torquato [15] have devised a method to produce large disordered hyperuniform particle dispersions based on spatial tessellations of point patterns, but their procedure is distinct from what we consider herein.
@@
These works lack a direct characterization of the large-scale density fluctuations of the structures themselves derived from hyperuniform point patterns. In this work, for the first time, we characterize the degree to which the hyperuniformity of the point pattern is inherited by network structures composed of the edges of tiles from spatial tessellations of 2D and 3D hyperuniform point patterns (see Sec. II for more details). Specifically, we build on the configuration-to-network-to-3D printed structure pipeline introduced in Obrero et al. [17] for designing and manufacturing disordered network metamaterials. First, we generate uniformly randomized lattice (URL) [30] and disordered stealthy hyperuniform point patterns with varying degrees of short-scale translational disorder, as well as totally uncorrelated point patterns in $d=2,3$ under periodic boundary conditions. Then, we generate the Voronoi (V), Delaunay (D), Delaunay-Centroidal (C), and Gabriel (G) tessellations of these point patterns and use the edges of the tiles in these tessellations to form spatially embedded network structures (see also [24]).
@@
Subsequently, we characterize these structures in two ways. First, to characterize the structures of two-phase media derived from 2D networks across length scales, we apply a “thickness” to the network edges such that the network fills a non-zero fraction of the space, then compute the spectral density $\tilde{\chi}_{{}_{V}}(\mathbf{k})$. Second, inspired by traditional variance-based order characterization methods [31, 32, 33], we propose the use of a variance measure, the edge-length variance $\sigma^{2}_{\ell}(R)$, associated with the total edge length of a spatially embedded network within a spherical observation window of radius $R$, which we apply to our 2D and 3D networks. Such a variance measure allows us to analyze the structure of the networks themselves as opposed to two-phase structures derived from them, avoiding the need to make a particular choice for the edge thickness or shape when analyzing the structures.
@@
We find that the hyperuniformity of point patterns is not completely inherited by network structures derived from their spatial tessellations, and the degree to which hyperuniformity is inherited varies across different point patterns, tessellation types, and space dimensions. Such subtleties in the inheritance of hyperuniformity by the network structures are consistent with the varying rank-orderings of physical properties for the 2D networks derived from different tessellations of hyperuniform and nonhyperuniform point patterns in Ref. 14. More specifically, for 2D networks derived from disordered hyperuniform point patterns with small degrees of local translational order, we find that G networks do not inherit the hyperuniformity of the progenitor point patterns, while D, V, and C are only effectively hyperuniform with C networks inheriting hyperuniformity most effectively. Furthermore, increasing the local translational disorder does not change the hyperuniformity of the underlying point pattern, but it does increase the values of the hyperuniformity index $H$ for the D, V, and C networks. For large degrees of local translational disorder, we find all the networks are nonhyperuniform.
@@
We also demonstrate that our proposed measure $\sigma^{2}_{\ell}(R)$ has analogous scaling behaviors to the number variance $\sigma_{N}^{2}(R)$, and detects the same structural characteristics at intermediate- and large-length scales in our network structures as the well-established $\tilde{\chi}_{{}_{V}}(k)$ calculation. This indicates that the incomplete inheritance of hyperuniformity by the two-phase network media is intrinsic to the tessellation-based network structures and not a result of a particular choice of edge thickness (or shape, for $d=3$) when mapping the network structures to two-phase media. Having established the correspondence between traditional variance measures and $\sigma_{\ell}^{2}(R)$ for the 2D networks, we apply the same method to 3D networks and show that V networks derived from disordered hyperuniform point patterns tend to have the greatest suppression of large-scale density fluctuations, followed by C, D, and G networks. This change in optimal network type from C networks for $d=2$ to V networks for $d=3$ is not unexpected given that the currently known structures that minimize large-scale density fluctuations also change across space dimensions [18].
@@
The rest of the paper is organized as follows. Section II contains pertinent mathematical background for the methods used to generate and characterize the point patterns and networks used in this work. In Sec. III, we discuss how hyperuniformity can be extended to spatial networks and introduce the edge-length variance $\sigma_{\ell}^{2}(R)$. In Sec. IV, we characterize the structures of these networks using the spectral density and edge-length variance. In Sec. V we offer concluding remarks and outlook for future studies.
@@ II Mathematical Background and Methods
In this section, we review the quantification of the pair statistics of point patterns and two-phase heterogeneous media and how they can be used to determine and characterize the degree of hyperuniformity in a given system. Then, we introduce the one nonhyperuniform and two hyperuniform point pattern models considered herein. Finally, we describe the four tessellation schemes we use to convert these point patterns into networks and the method used to convert these networks into two-phase media.
@@ II.1 Pair statistics
$S(\mathbf{k})=1+\rho\tilde{h}(\mathbf{k}),$ (1)
@@
$\tilde{\chi}_{{}_{V}}(\mathbf{k})=\frac{|\sum_{j=1}^{N_{p}}e^{-i\mathbf{k}\cdot{\mathbf{r}_{j}}}\tilde{m}(\mathbf{k};A)|^{2}}{V},\;(\mathbf{k}\neq 0),$ (6)
@@ II.3 Hyperuniform and nonhyperuniform model point patterns
In this work, we consider one nonhyperuniform model and two hyperuniform model point patterns in $d=2,3$. The nonhyperuniform model we consider is a totally uncorrelated point process in $\mathbb{R}^{d}$. We generate these point patterns at a number density $\rho=1$ in a fixed $d$-dimensional box by uniformly randomly placing $N$ points independent of each other. Such a system with no spatial correlations has $g_{2}(r)=S(k)=1$ and thus is nonhyperuniform (i.e., $\alpha=0$) with $\sigma_{N}^{2}\sim R^{d}$ at large $R$.
One hyperuniform model we consider is the uniformly randomized lattice (URL) [30]. Here, we make the particular choice of displacing the points of the $\mathbb{Z}^{d}$ lattice by a random $d$-dimensional vector uniformly distributed on a scaled version of the unit cell of $\mathbb{Z}^{d}$, i.e., $a\mathcal{F}\equiv[-a/2,a/2)^{d}$, where $a$ controls the perturbation strength. Regardless of the exact value of $a>0$, the URL is a class I hyperuniform system with $\alpha=2$. One can tune the degree of local translational disorder in the point pattern with the specific choice of $a$; in particular, larger values of $a$ result in a greater degree of local translational disorder. Here, we consider values of $a$ between 0.1 and 1.75.
Additionally, we consider disordered stealthy hyperuniform point patterns, which have $S(k)=0$ for $0<k\leq K$ and thus are class I hyperuniform. We generate these point patterns using the collective coordinate optimization scheme starting from totally uncorrelated initial conditions [37]. The parameter $\chi$ is a dimensionless measure of the ratio of constrained degrees of freedom in the stealthy system (i.e., the number of wavenumbers less than $K$ constrained to be 0) to the total number of degrees of freedom (roughly $dN$, where $N$ is the number of particles in the system). Relatively unconstrained (low-$\chi$) disordered stealthy systems will have greater translational disorder on short length scales and as $\chi$ increases within the so-called “disordered regime” ($20<\chi<1/$ 2 for $d=2,3$ [19]) the degree of short length scale translational disorder decreases. In this work, we consider $\chi=0.2,0.4,$ and 0.48.
@@ II.4 Network generation
To generate the networks examined herein, we first map point patterns under periodic boundary conditions in $\mathbb{R}^{2}$ ($\mathbb{R}^{3}$) to sets of polygons (polyhedra) that tile the space using the Voronoi (V), Delaunay (D), Delaunay-centroidal (C), and Gabriel (G) tessellations (described below). Then, we produce a network from the edges that make up the boundary of each cell in the tessellation. Each polygonal cell in the Voronoi tessellation is the region of space closer to one point than to any other point in the progenitor point pattern. The Delaunay tessellation is the dual network of the Voronoi tessellation. The Delaunay-centroidal tessellation [9] is generated by connecting the centroid of each simplex in the Delaunay tesselation to the centroid of each neighboring simplex, i.e., those that share an edge (or face in $\mathbb{R}^{3}$). Finally, the Gabriel tessellation [38] is derived from the Delaunay tessellation by removing any edge if a sphere generated using that edge as its diameter contains another vertex in the tessellation.
@@
In addition to considering these sets of edges, in $\mathbb{R}^{2}$ we consider the two-phase medium produced when a rectangle of dimensions $2\delta\times b+2\delta$ is superposed on each edge, where $b$ is the Euclidean length of the original edge and $\delta$ is some prescribed amount by which the edge is “thickened.” Here, we choose $\delta$ to be equal to one twentieth of $\rho^{-1/d}$. We note here that the particular choice of beam thickness only impacts the small-scale structural characteristics of the medium; the large-scale density fluctuations are not significantly impacted. In Fig. 1 we show an example of each of these types of networks. Because this mapping of point patterns to two-phase media is a perturbation of the underlying network structure, we additionally desire a method to directly examine the network structure. We introduce and discuss the merits of such a method in Sec. III
III Toward a Generalization of Hyperuniformity to Spatial Networks
In the previous section, we described how we can decorate network structures embedded in $\mathbb{R}^{2}$ to generate two-phase media, so that we can use standard techniques (e.g., the spectral density or volume fraction variance) to characterize their density fluctuations across length scales. However, treatment of network structures as a set of edges (i.e., without decorations) embedded in $\mathbb{R}^{d}$ are objects of great interest and have wide-ranging applications from the modeling of public transportation systems to neuron activity in the brain, among many others (see, e.g., Ref. 39 and references therein). Moreover, for the purposes of the present work, it is not yet known how the decoration scheme chosen here perturbs the density fluctuations of the underlying network structure, so we desire a way to probe the structural characteristics of the network directly.
To carry out this characterization, we propose computing the edge-length variance $\sigma^{2}_{\ell}(R)$, which is the variance of the total edge length of a network embedded in $\mathbb{R}^{d}$ within a uniformly randomly positioned observation window of radius $R$. Such a characterization method is motivated by the collection of variance measurements that were recently shown to effectively characterize the translational disorder in point patterns [33], two-phase media [31], and scalar fields [32] across length scales. For each (hyper)spherical sampling window placed into the system, we sum all of the Euclidean lengths of edges that fall entirely within the window. For edges that fall partially within the window (i.e., one vertex of the edge is inside the window and the other is outside), we consider only the length of the portion of the edge that lies within the window; such a treatment of edge length is naturally connected to the concept of metric networks [40]. While in this work the edges in our networks are weighted by their Euclidean length, this method can be generalized to treat spatially embedded networks with arbitrary weighting schemes. In these more general cases, we instead sum the weights of all of the edges that lie entirely within the window, and for edges that fall partially within the window we add a fraction of the weight of that edge proportional to the fraction of the Euclidean edge length that falls within the sampling window.
For the two-phase media generated using the scheme in Sec. IID, the edge-length variance can be linked to the volume-fraction variance by first noting that the volume fraction $\phi$ of the network edges with a prescribed thickness (see previous section for the 2D construction) in an observation window with radius $R$ is approximately proportional to the total length $\ell$ of the edges in the window. Additionally, there is a correction term that accounts for the overlap of the thickened edges. Together, the relationship between volume fraction $\phi$ and total edge length $\ell$ is given by
$\phi=\frac{\ell\delta+O\big{(}\delta^{d}N(R)\big{)}}{v_{1}(R)}$ (13)
in the limit that the edge thickness $\delta\rightarrow 0$, where $N(R)$ is the number of network nodes within the spherical region with radius $R$. Thus, the volume fraction variance $\sigma^{2}_{V}(R)$ is linked to $\sigma_{\ell}^{2}(R)$ via
$\sigma_{V}^{2}(R)=\frac{\delta^{2}\sigma^{2}_{\ell}(R)+O\big{(}\delta^{2d}\sigma_{N}^{2}(R)\big{)}}{v_{1}^{2}(R)}\quad\textrm{as }\delta\rightarrow 0.$ (14)
From this, we can see that the large-$R$ scalings of $\sigma_{V}^{2}(R)$ given in Eq. (II.2) correspond to the edge-length variance scalings
$\sigma_{\ell}^{2}(R)\sim\begin{cases}R^{d-1},\quad\alpha>1&\textrm{(class I)}\\ R^{d-1}\textrm{ln}(R),\quad\alpha=1&\textrm{(class II)}\\ R^{d-\alpha},\quad 0<\alpha<1&\textrm{(class III)},\end{cases}$ (15)
for hyperuniformity and $R^{d}$ for nonhyperuniformity. We note that such scalings correspond to those of the number variance $\sigma_{N}^{2}(R)$ for point patterns given in Eq. (9). Deriving the corresponding scattering function for these network structures to supplement $\sigma_{\ell}^{2}(R)$ is an important open problem in the field.
While the spectral density methods used in this work are reasonably fast—on the order of several minutes per structure for the system sizes used herein when run on a Windows 11 operating system with 16 GB of RAM and a 12th Gen Intel Core i7-1255U processor running at 1.70 GHz—they suffer from potential discretization errors, especially when generating the pixelizations of very large or nearly crystalline networks. Moreover, one must choose a particular beam thickness (and shape for $d>2$), which may impact the density fluctuations of the structure in unexpected ways. Both of the aforementioned issues can be avoided when computing $\sigma^{2}_{\ell}(R)$ because the intersections between line segments (the network edges) and hyperspheres (the observation windows of radius $R$) can straightforwardly be computed exactly and no decoration of the network structure is required to do the computations. The volume-fraction variance can also be computed for some two-phase media without discretization, but computing the overlap of an observation window with such a medium is more computationally intensive than the corresponding overlap calculation for $\sigma^{2}_{\ell}(R)$. Computing $\sigma^{2}_{\ell}(R)$ for a network structure derived from a $N=300^{2}$ 2D point pattern takes on the order of an hour, while the corresponding volume fraction variance calculation for the overlapping rectangle system takes several hours, where both calculations are done using the same computer specifications as above.
IV Results
IV.1 Two-dimensional networks